Brief Exercise 9-56
Bonds Issued at a Discount (Effective Interest)

Crafty Corporation received $472,088 of cash upon issuance of 500 $1,000 par value bonds. Each bond has a stated rate of 5% and will mature on December 31, 2026, 7 years after the issuance of the bonds. Interest is paid annually on December 31. The market rate of interest is 6%.

Required:

Prepare the journal entry for December 31, 2022. If required, round amounts to the nearest whole dollar.

2022 Dec. 31 debit Credit
(Record interest expense.)

Prepare the journal entry for December 31, 2023. If required, round amounts to the nearest whole dollar.

2023 Dec. 31 debit credit
(Record interest expense.)
Answer #1

Interest expense 2020 = 472088*6% = 28325-25000 = 3325

Interest expense 2021 = (472088+3325)*6% = 28525-25000 = 3525

Prepare the journal entry for December 31, 2022. If required, round amounts to the nearest whole dollar.

2022 Dec. 31 Interest expense (478938*6%) 28736
Discount on bonds payable 3736
Bonds payable 25000
(Record interest expense.)

Prepare the journal entry for December 31, 2023. If required, round amounts to the nearest whole dollar.

2023 Dec. 31 Interest expense (482674*6%) 28960
Discount on bonds payable 3960
Cash 25000
(Record interest expense.)